Is the point (-4 1/2) on the graph of y=1/2x-4? How do you know?
Found 2 solutions by stanbon, god2012:
Answer by stanbon(75887)

Is the point (-4, 1/2) on the graph of y=1/2x-4?
How do you know?
----
Test it.
1/2 = (1/2)(-4) -4
---
1/2 = -2-4
---
1/2 = -6
---
False
---
That point is not on that line.
Cheers,
Stan H.

Answer by god2012(113)

Point P(-4, 1/2) and y = 1/2x - 4
Put x = -4 in the equation
y = 1/2 * (-4) -4 = -6
But y = 1/2 in point p
Therefore the point P is not on the graph